1.List some characteristics of fluid pressureand SHM.
2.State first law of thermodynamics and show for different thermodynamic system.
3.When you walk along a road and a track passes you at high speed. Why do you feel pulled towards the truck?
4.A system absorbs 35J of heat and in the process it does 11J of work. If the initial internal energy is 205J what is the final internal energy?
1
Expert's answer
2020-02-03T14:16:38-0500

1. The properties of fluid pressure are:

•The pressure in a fluid increases with depth. The pressure of water is highest at the lowest point of the cylinder as the water spurts out the furthest from this point.

•The pressure in a fluid does not depend on the shape or size of the container.

•The pressure in a liquids acts in all direction.

2. The first law of thermodynamics is the application of the conservation of energy principle to heat and thermodynamic processes (ΔU=Q+W) .

The first law of thermodynamics simply states that energy can be neither created nor destroyed. Thus power generation processes and energy sources actually involve conversion of energy from one form to another, rather than creation of energy from nothing.

3. A space conflict happens when two road users try to move into the same space at the same time. To drive safely, you need to keep areas of space — called space margins — around your vehicle. If you need to stop suddenly, a driver too close behind you could cause a space conflict.

4. Here, q = +35 J, w = – 11 J

 According to first law of thermodynamics,

U = q + w = 35 + (–11) = 24 J

Thus, internal energy of the system increases by 10 J.